# Understanding Before Recommendation: Semantic Aspect-Aware Review Exploitation via Large Language Models

### v1

@@ Abstract.
Recommendation systems harness user-item interactions like clicks and reviews to learn their representations. Previous studies improve recommendation accuracy and interpretability by modeling user preferences across various aspects and intents. However, the aspects and intents are inferred directly from user reviews or behavior patterns, suffering from the data noise and the data sparsity problem. Furthermore, it is difficult to understand the reasons behind recommendations due to the challenges of interpreting implicit aspects and intents. Inspired by the deep semantic understanding offered by large language models (LLMs), we introduce a chain-based prompting approach to uncover semantic aspect-aware interactions, which provide clearer insights into user behaviors at a fine-grained semantic level. To incorporate the abundant interactions of various aspects, we propose the simple yet effective Semantic Aspect-based Graph Convolution Network (short for SAGCN). By performing graph convolutions on multiple semantic aspect graphs, SAGCN efficiently combines embeddings across multiple semantic aspects for final user and item representations. The effectiveness of the SAGCN was evaluated on three publicly available datasets through extensive experiments, which revealed that it outperforms all other competitors. Furthermore, interpretability analysis experiments were conducted to demonstrate the interpretability of incorporating semantic aspects into the model.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommendation systems have become essential techniques for various online platforms. It not only assists users in navigating through overwhelming volumes of information but also serves as a powerful tool for service providers to increase customer retention and revenue. Within this field, Collaborative Filtering (CF)-based models have achieved significant advancements in learning the user and item representations by effectively exploiting the historical user-item interactions (Bell and Koren, 2007; He et al., 2017; Wu et al., 2016; Xue et al., 2017). For instance, Matrix Factorization (MF) decomposes the user-item interaction matrices into lower-dimensional latent vectors and predicts the user preference with the inner product (Bell and Koren, 2007). By replacing the inner product with nonlinear neural networks, NeuMF (He et al., 2017) can model the complex interaction behaviors, yielding better recommendation performance. More recently, Graph Convolution Networks (GCNs)-based models (van den Berg et al., 2018; Wang et al., 2019b, a; He et al., 2020; Wang et al., 2020a; Liu et al., 2021; Fan et al., 2022) offer a cutting-edge approach in recommendation systems by leveraging the high-order connectivities based on the topological structure of the interaction data. For example, NGCF (Wang et al., 2019b) incorporates high-order connectivity of user-item interactions in the user-item bipartite graph for representation learning in recommendation.
[…]
@@
Semantic aspects often denote specific features or attributes of a product or service in which users might be interested in the context of recommendation systems and sentiment analysis. Extracting semantic aspects from reviews could enhance our comprehension of the complex relations between each user and item at a finer level. Moreover, semantic aspects, like meaningful attributes or concepts of an item, provide a deeper understanding of its inherent features, thereby enhancing the interpretability of recommendations. As shown in Fig. 1, the review might mention various semantic aspects such as functionality, durability, and ease of use. Each of these aspects carries a specific meaning related to user experience. By analyzing and summarizing the raw review into structured semantic aspects, we can better understand the underlying rationales behind user behaviors. Focusing on reviews that highlight specific item aspects enables us to discern semantic aspect-aware interactions. For instance, a review might emphasize aspects like functionality and durability but omit ease of use. Such patterns indicate distinct interactions between the user and item concerning functionality and durability, but not for ease of use. Leveraging these semantic-aware insights leads to richer and more accurate representations of both users and items.
@@
Figure 1. An example of extracting semantic aspect-aware reviews (e.g., functionality, durability, and ease of use) from user review via LLM. Based on the extracted reviews, the semantic aspect-aware interactions are discerned.
@@
Large Language Models (LLMs) have emerged as powerful tools in natural language processing and have recently gained significant attention in various fields (Da’u and Salim, 2022; Sileo et al., 2022). Their contextual understanding allows for accurately distilling various semantic aspects based on reviews without being affected by noise present in reviews. Moreover, based on the extracted aspects, we can form aspect-specific interactions, capturing users’ preferences from various semantic aspects. Motivated by the above consideration, we devise a chain-based prompting approach, which consists of the semantic aspects extraction prompt and semantic aspects-aware review extraction prompt. The former prompt extracts semantic aspects from raw reviews, while the latter one identifies reviews aligned with these distinct semantic aspects. To demonstrate the potential of structuring raw reviews into semantic aspects via LLMs in recommendations, we introduce a Semantic Aspect-based Graph Convolution Network (SAGCN) model. This model performs graph convolution operations on aspect-aware interaction graphs, each of which is constructed based on the user-item interactions of a specific aspect derived from the aspect-aware reviews. The final user and item representations are obtained by combining the embeddings learned from different aspect-specific graphs. The merits of our model are twofold: (1) embeddings are enriched by learning from various aspects; (2) interpretability is improved by incorporating semantic aspects. Extensive experiments are conducted on three benchmark datasets to evaluate the SAGCN. Results show that SAGCN can gain better user and item representations than SOTA GCN-based recommendation models. The interpretability analysis experiments demonstrate the interpretability of incorporating semantic aspects into our model.
@@
  • We emphasize the significance of utilizing semantic aspect-aware interaction in recommendations. Inspired by this, we propose a chain-based prompting approach to extract the semantic aspect-aware user-item reviews, which are used to discern aspect-aware interactions.
  • We introduce an SAGCN model to leverage the extracted semantic aspects via LLM, demonstrate the effectiveness of the extracted semantic aspects by our method on enhancing the recommendation accuracy and interpretability.
  • We conducted empirical studies on three benchmark datasets to evaluate the performance of SAGCN. Our results demonstrate that SAGCN outperforms state-of-the-art GCN-based recommendation models in terms of accuracy and interpretability.
@@
2.1. Collaborative Filtering
@@
Collaborative Filtering (CF) techniques (Koren et al., 2009; Rendle et al., 2009; Wang et al., 2019a; He et al., 2017; Wang et al., 2019b; He et al., 2020; Liu et al., 2019) play an crucial role in recommendation systems. The core idea is to learn users’ and items’ representations ($aka.$ embedding vectors) by exploiting the collaborative information based on user-item interactions. With the learned user and item embedding vectors, an interaction function is adopted to predict a user’s preference for an item. For example, Matrix Factorization (MF) (Koren et al., 2009) learns the user and item embeddings by minimizing the error of re-constructing the user-item interaction matrix. Later, due to its powerful capability, the deep learning technique is widely adopted in recommendations to learn better user and item embeddings (Xue et al., 2017) or model more complicated interactions between users and items (He et al., 2017; Li et al., 2019). For example, NeuMF (He et al., 2017) leverages a nonlinear neural network to model the complex interactions between users and items. Owing to the powerful ability to learn representations based on the graph structure, graph convolution networks are employed to exploit the high-order user-item interactions for improving recommendation performance (van den Berg et al., 2018; Wang et al., 2019b). A typical method is NGCF (Wang et al., 2019b), which aggregates embeddings from high-order neighbors by performing graph convolution on the user-item interaction graph.
[…]
@@ 2.2. GCN-based Recommendation Models
In recent years, Graph Convolutional Networks (GCNs) have garnered significant attention in the realm of recommendation systems, owing to their robust capabilities in learning node representations from non-Euclidean structures (van den Berg et al., 2018; Wang et al., 2019b; Ying et al., 2018; Wang et al., 2019a, 2020a; He et al., 2020; Liu et al., 2021; Lin et al., 2022; Liu et al., 2022b; Cheng et al., 2023). This also led to the rapid advancement of GCN-based recommendation models. For example, GC-MC (van den Berg et al., 2018) employs one convolution layer to exploit the direct connections between users and items. NGCF (Wang et al., 2019b) propagates embeddings on the user-item interaction graph to exploit the high-order proximity. Inspired by the study of simplifying GCN (Wu et al., 2019), researchers also introspect the complex design in GCN-based recommendation models. For example, He et al. (He et al., 2020) observed that the two common designs feature transformation and nonlinear activation have no positive effects on the final performance. Thus, they proposed LightGCN, which removes these two modules and significantly improves recommendation performance.
To achieve robust node representation, fine-grained interactions are exploited to refine the graph structure. Wang et al. (Wang et al., 2020a) presented a GCN-based model, which yields disentangled representations by modeling a distribution over intents for each user-item interaction; IMP-GCN (Liu et al., 2021) learns user and item embeddings by performing high-order graph convolution inside subgraphs constructed by the user nodes with common interests and associated items. By incorporating side information, GRCN (Wei et al., 2020) adjusts the structure of the interaction graph adaptively based on the status of model training for discovering and pruning potential false-positive edges; and LATTICE (Zhang et al., 2021) mines the latent structure between items by learning an item-item graph. Graph convolutional operations are performed on both item-item and user-item interaction graphs to learn user and item representations.
2.3. LLMs in Recommendation
Large Language Models (LLMs) have emerged as powerful tools in natural language processing and have recently gained increasing attention in recommendation systems. The LLM-based models are particularly proficient at capturing contextual information and comprehending user queries, item descriptions, and other textual data in a highly effective manner (Geng et al., 2022; Da’u and Salim, 2022; Sileo et al., 2022; Wu et al., 2023). By understanding the context, LLM-based models can improve the recommendation performance.
There are two main categories for incorporating LLMs into recommendation systems. One is to adapt LLMs as recommendation models. Most generative model-based work translates recommendation tasks as natural language tasks and then applies techniques such as in-context learning, prompt tuning, and instruction tuning to adapt LLMs to generate the recommendation results (Ji et al., 2023; Zhang et al., 2023b, a; Kang et al., 2023; Li et al., 2023). Moreover, with the impressive capabilities demonstrated by ChatGPT, this type of work has received increased attention recently. Ji et al. (Ji et al., 2023) proposed a generative recommendation method that uses LLMs to generate recommended items. The other is to extract high-quality representations of textual features or leverage the extensive external knowledge encoded within them (Liu et al., 2023). Most existing works align the representations of pre-trained models like BERT (Devlin et al., 2019) with the domain-specific data through fine-tuning. In addition to taking LLMs as recommendation systems, some studies also utilize LLMs to construct model features. For example, GENRE (Liu et al., 2023) introduced three prompts to conduct three feature enhancement sub-tasks for news recommendation. The prompts involved refining news titles using ChatGPT, extracting profile keywords from user reading history, and generating synthetic news to enhance user interactions based on their history. By incorporating these LLM-constructed features, traditional news recommendation models can be greatly improved.

### v2

@@ Abstract.
Recommendation systems harness user-item interactions like clicks and reviews to learn their representations. Previous studies improve recommendation accuracy and interpretability by modeling user preferences across various aspects and intents. However, the aspects and intents are inferred directly from user reviews or behavior patterns, suffering from the data noise and the data sparsity problem. Furthermore, it is difficult to understand the reasons behind recommendations due to the challenges of interpreting implicit aspects and intents. To address these constraints, we harness the sentiment analysis capabilities of Large Language Models (LLMs) to enhance the accuracy and interpretability of the conventional recommendation methods. Specifically, inspired by the deep semantic understanding offered by LLMs, we introduce a chain-based prompting strategy to uncover semantic aspect-aware interactions, which provide clearer insights into user behaviors at a fine-grained semantic level. To incorporate the rich interactions of various aspects, we propose the simple yet effective Semantic Aspect-based Graph Convolution Network (short for SAGCN). By performing graph convolutions on multiple semantic aspect graphs, SAGCN efficiently combines embeddings across multiple semantic aspects for final user and item representations. The effectiveness of the SAGCN was evaluated on four publicly available datasets through extensive experiments, which revealed that it outperforms all other competitors. Furthermore, interpretability analysis experiments were conducted to demonstrate the interpretability of incorporating semantic aspects into the model.
@@
ccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommendation systems have become essential components of various online platforms. They not only assist users in navigating through overwhelming volumes of information but also serve as a powerful tool for service providers to increase customer retention and revenue. Within this field, Collaborative Filtering (CF)-based models have achieved significant advancements in learning the user and item representations by effectively exploiting the historical user-item interactions (Bell and Koren, 2007; He et al., 2017; Wu et al., 2016; Xue et al., 2017). For instance, Matrix Factorization (MF) decomposes the user-item interaction matrices into lower-dimensional latent vectors and predicts the user preference with the inner product (Bell and Koren, 2007). By replacing the inner product with nonlinear neural networks, NeuMF (He et al., 2017) can model the complex interaction behaviors, resulting in better recommendation performance. More recently, Graph Convolution Networks (GCNs)-based models (van den Berg et al., 2018; Wang et al., 2019b, a; He et al., 2020; Wang et al., 2020a; Liu et al., 2021; Fan et al., 2022) offer a cutting-edge approach in recommendation systems by leveraging the high-order connectivities based on the topological structure of the interaction data. For example, NGCF (Wang et al., 2019b) incorporates high-order connectivities between user and item nodes in the bipartite graph to learn their representations for recommendation.
To better understand user preferences, researchers have paid significant attention to exploiting fine-grained behaviors between users and items on different aspects or intents (Cheng et al., 2018b, a; Chin et al., 2018; Liu et al., 2019; Chen et al., 2024). Existing methods can be categorized into two main approaches. First, employing topic models (Blei et al., 2003), the statistical model designed to identify the abstract “topics” based on the frequency and co-occurrence of words in textual content, to extract aspects from these reviews. These extracted aspects are then incorporated into user preference modeling (Cheng et al., 2018b, a). In particular, ALFM (Cheng et al., 2018b) applies a topic model on reviews to extract the topics the user pays attention to, while A3NCF (Cheng et al., 2018a) learns the user attention from reviews via a neural attention network in an end-to-end manner. While these models can identify topics or aspects in the review, understanding their semantic meaning is not straightforward. In addition, the noise in the review has a negative impact on the aspect extraction. For example, words that have no contribution to user preference, like “the”, “last”, and “year” in reviews, are extracted by the topic models and leveraged into the training process of the recommendation model. Second, disentangling the underlying aspect/factor behind interaction data (Wang et al., 2020b, a; Liu et al., 2023a; Li et al., 2024). Disentangled representations are becoming increasingly important for learning robust and independent representations from user-item interaction data in recommendation (Wang et al., 2020b, a). For example, DisenHAN (Wang et al., 2020b) learns disentangled user/item representations from different aspects in a heterogeneous information network, while DGCF (Wang et al., 2020a) models a distribution over intents for each user-item interaction to study the diversity of user intents on adopting the items. Despite their ability to accurately model user preference and capture the underlying factors, such models extract the aspects relying on user-item interaction data, which suffer from data sparsity and bias issues in recommendations. In addition, the implicit aspects and intents behind the recommendations are often elusive and not readily interpretable.
@@
Semantic aspects often denote specific features or attributes of a product or service in which users might be interested in the context of recommendation systems and sentiment analysis. Extracting semantic aspects from reviews could improve our comprehension of the complex relations between each user and item at a finer level. Moreover, semantic aspects, like meaningful attributes or concepts of an item, provide a deeper understanding of its inherent features, thereby enhancing the interpretability of recommendations. As shown in Fig. 1, the review might mention various semantic aspects such as functionality, durability, and ease of use. Each of these aspects carries a specific meaning related to user experience. By analyzing and summarizing the raw review into structured semantic aspects, we can better understand the underlying rationales behind user behaviors. Focusing on reviews that highlight specific item aspects enables us to discern semantic aspect-aware interactions. For instance, a review might emphasize aspects like functionality and durability but omit ease of use. Such patterns indicate distinct interactions between the user and item concerning functionality and durability, but not for ease of use. Utilizing these semantic-aware insights leads to richer and more accurate representations of both users and items.
@@
Figure 1. An example of extracting semantic aspect-aware reviews (e.g., functionality, durability, and ease of use) from user review via LLM. Based on the extracted reviews, the semantic aspect-aware interactions are discerned. The results from LLM indicate that user $u$ interacts with item $i$ in terms of functionality and durability aspects, but dose not interacts it in relation to ease of use.
@@
Large Language Models (LLMs) have emerged as powerful tools in natural language processing and have recently gained significant attention in various fields (Da’u and Salim, 2022; Sileo et al., 2022). Their contextual understanding allows for distilling various semantic aspects based on reviews without being affected by noise present in reviews. In this paper, we propose a novel approach that combines the strengths of LLMs with conventional recommendation techniques by extracting the semantic aspects from user reviews. Although LLMs exhibit advanced proficiency across numerous tasks, they still encounter challenges in accurately identifying aspects, especially in autonomously discovering unseen aspects without explicit instructions. To mitigate this constraint of LLMs, we devise a chain-based prompting strategy, which consists of the semantic aspects extraction prompt and semantic aspects-aware review extraction prompt. The first prompt extracts semantic aspects from raw reviews, after which we refine and select high-quality aspects from the entire dataset. The second prompt then identifies reviews that align with these high-quality semantic aspects. On the basis of the extracted aspect-aware reviews, we can form aspect-specific interactions, capturing users’ preferences from various semantic aspects. To demonstrate the potential of structuring raw reviews into semantic aspects via LLMs in recommendations, we present a Semantic Aspect-based Graph Convolution Network (SAGCN) model. This model performs graph convolution operations on aspect-aware interaction graphs, each of which is constructed based on the user-item interactions of a specific aspect derived from the aspect-aware reviews. The final user and item representations are obtained by combining the embeddings learned from different aspect-specific graphs. The merits of our model are twofold: (1) embeddings are enriched by learning from various aspects; (2) interpretability is improved by incorporating semantic aspects. Extensive experiments have been conducted on three benchmark datasets to evaluate the SAGCN. The results indicate that SAGCN outperforms state-of-the-art GCN-based recommendation models in generating more effective user and item representations. Additionally, interpretability analysis experiments demonstrate the effectiveness of integrating semantic aspects into our model. To facilitate further research, we have released the code and the constructed semantic aspect-aware interaction data for others to replicate this work  [^1] https://github.com/HuilinChenJN/LLMSAGCN..
@@
  • We emphasize the significance of utilizing semantic aspect-aware interaction in recommendations. Inspired by this, we propose a novel approach that combines the strengths of LLMs with conventional recommendation techniques by extracting the semantic aspects from user reviews.
  • To extract the semantic aspects and semantic aspect-aware reviews, we devise a chain-based prompting strategy. On the basis of this, we could discern aspect-aware interactions within each user-item review.
  • To demonstrate the effectiveness of our approach in enhancing recommendation accuracy and interoperability, we present the SAGCN model that leverages the extracted semantic aspect-aware interactions for user preference modeling.
  • We conducted empirical studies on four benchmark datasets to evaluate the performance of SAGCN. Our results demonstrate that SAGCN outperforms state-of-the-art recommendation models in both accuracy and interpretability.
@@
This section briefly introduces recent advances in collaborative filtering based on user reviews, GCN-based recommendation models, and large language models in recommendation, which are closely relevant to our work.
2.1. Collaborative Filtering Based on User Reviews
@@
Collaborative Filtering (CF) techniques (Koren et al., 2009; Rendle et al., 2009; Wang et al., 2019a; He et al., 2017; Wang et al., 2019b; He et al., 2020; Liu et al., 2019) play a pivotal role in recommendation systems. The fundamental concept involves learning representations (or embedding vectors) of users and items by harnessing collaborative information derived from user-item interactions. These learned user and item embedding vectors are then utilized within an interaction function to predict a user’s preference for an item. For example, Matrix Factorization (MF) (Koren et al., 2009) learns the user and item embeddings by minimizing the error of re-constructing the user-item interaction matrix. Due to the powerful capabilities of deep learning, it is widely adopted in recommendations to improve the quality of user and item embeddings (Xue et al., 2017) or model more complicated interactions between users and items (He et al., 2017; Li et al., 2019). One prominent method is NeuMF (He et al., 2017), which leverages a nonlinear neural network to model complex interactions between users and items.
To mitigate the data sparsity problem, user reviews (Catherine and Cohen, 2017; McAuley and Leskovec, 2013; Tan et al., 2016; Cheng et al., 2018a; Chen et al., 2018; Cheng et al., 2018b; Chin et al., 2018; Wei et al., 2023; Liu et al., 2023b) have been leveraged in recommender systems to learn the user and item representations as they can provide additional information about user preference. One of the typical methods to learn user preferences by exploiting reviews is through deep learning techniques (Zheng et al., 2017). For example, DeepCoNN (Zheng et al., 2017) learn item characteristics and user preferences jointly from user reviews via the deep model. Existing methods extract aspects from reviews and integrate these aspects into user preference modeling. For example, A3NCF (Cheng et al., 2018a) extracts user preferences and item characteristics from review texts via topic model to enhance the user’s and item’s representations. SULM (Bauman et al., 2017) uses a sentiment analysis system, Opinion Parser, to extract aspects. Through a matrix factorization approach, it calculates the personal impact each aspect has on the user’s overall rating. Furthermore, it is well-recognized that multimodal features can provide complementary information to each other, as demonstrated in (Zhang et al., 2016, 2017; Bauman et al., 2017; Liu et al., 2019). JRL (Zhang et al., 2017) first utilizes deep neural networks to separately extract user and item features from multiple modalities, such as ratings, reviews, and images, then forms the final user and item representations by concatenating those multimodal features. To model users’ diverse preferences, MAML (Liu et al., 2019) fuses user reviews with item images to estimate the weights for different aspects of items.
Despite their success, the above-mentioned methods are still incapable of computing the optimal embeddings for recommendation. This deficiency arises from the fact that collaborative filtering signals are only implicitly captured, thereby neglecting the transitivity property of behavioral similarity (Wang et al., 2019b). A common solution is to explicitly incorporate the graph structure to enhance user and item representation learning. In the next, we mainly discussed the Graph-based recommendation methods.
2.2. Large Language Models in Recommendation
Large Language Models (LLMs) have emerged as potent tools in natural language processing, garnering increasing attention in recommendation systems. The LLM-based models excel in capturing contextual information and comprehending user queries, item descriptions, and other textual data in a highly effective manner (Geng et al., 2022; Da’u and Salim, 2022; Sileo et al., 2022). By understanding the context, LLM-based models can improve the recommendation performance.
There are two primary approaches to integrating LLMs into recommendation systems. Firstly, LLMs can be adapted directly as recommendation models. Many works in this vein treat recommendation tasks as natural language tasks, employing techniques such as in-context learning, prompt tuning, and instruction tuning to tailor LLMs for generating recommendation results. (Ji et al., 2024; Zhang et al., 2023b, a; Kang et al., 2023; Li et al., 2023). Recent attention has been particularly drawn to this line of research, buoyed by the impressive capabilities showcased by models like ChatGPT. For instance, Ji et al. (Ji et al., 2024) proposed a generative recommendation method that uses LLMs to generate recommended items. PEPLER (Li et al., 2023) generates natural language explanations for recommendations by treating user and item IDs as prompts. Secondly, LLMs can be leveraged to extract high-quality representations of textual features or tap into the extensive external knowledge they encapsulate (Liu et al., 2024a). Most existing efforts align pre-trained models like BERT (Devlin et al., 2019) with domain-specific data through fine-tuning. Additionally, beyond serving as standalone recommendation systems, some studies employ LLMs to construct model features. For example, GENRE (Liu et al., 2024a) introduced three prompts to conduct three feature enhancement sub-tasks for news recommendation. The prompts involved refining news titles using ChatGPT, extracting profile keywords from user reading history, and generating synthetic news to enhance user interactions based on their history. By incorporating these LLM-constructed features, traditional news recommendation models can be greatly improved. RLMRec (Ren et al., 2024) generates user and item profiles based on interaction data and textual information, such as product descriptions, user reviews, and product titles. Then, these generated profiles are encoded as semantic information of LLMs and incorporated into the CF-based recommendation system.
Utilizing LLM’s sentiment analysis capability, we aim to better understand users’ various intentions expressed in reviews. Specifically, we first extract semantic aspects from user reviews and then identify the semantic-aspect interactions for each user and item pair. By leveraging semantic-aspect interactions, we can capture fine-grained user preferences. Compared to conventional methods that use features directly extracted from reviews by pretrained language models, semantic-aware interactions can improve the interpretability of recommendation models.
@@ 2.3. GCN-based Recommendation Models
Early recommendation approaches employ graph-based approaches to infer indirect user-item preferences. The high-order proximity between users and items is leveraged through random walks on the graph in order to provide effective recommendations (Gori and Pucci, 2007; Fouss et al., 2007). The recently proposed approaches exploit the user-item interaction graph to enrich the user-item interactions (Yang et al., 2018; Yu et al., 2018) as well as to explore additional collaborative relations, such as similarities between users and items (Chen et al., 2019; Yu et al., 2018). For example, WalkRanker (Yu et al., 2018) and CSE (Chen et al., 2019) utilize random walks to investigate the high-order proximity in both user-user and item-item relationships. These methods rely on random walks to sample interactions for model training, thus their performance is heavily depends on the quality of the generated interactions.
@@
In recent years, Graph Convolutional Networks (GCNs) have garnered significant attention in the realm of recommendation systems, owing to their robust capabilities in learning node representations from non-Euclidean structures (van den Berg et al., 2018; Wang et al., 2019b; Ying et al., 2018; Wang et al., 2019a, 2020a; He et al., 2020; Liu et al., 2021; Lin et al., 2022; Cheng et al., 2023; Ren et al., 2023; Liu et al., 2024b). This also led to the rapid advancement of GCN-based recommendation models. For instance, NGCF (Wang et al., 2019b) propagates embeddings on the historical interaction graph to exploit the high-order proximity. Some research has been dedicated to simplifying GCN (graph convolutional network) and GCN-based recommendation models (Wu et al., 2019). As a typical method, He et al. (He et al., 2020) observed that the two common designs feature transformation and nonlinear activation have no positive effects on the final performance. Hence, they proposed LightGCN, which removes these two modules and significantly improves recommendation performance. To obtain robust user and item representations, fine-grained interactions are exploited to refine the graph structure. Wang et al. (Wang et al., 2020a) presented a GCN-based model, which yields disentangled representations by modeling a distribution over intents for each user-item interaction; IMP-GCN (Liu et al., 2021) learns user and item embeddings by performing high-order graph convolution inside subgraphs constructed by the user nodes with common interests and associated items. By incorporating side information, GRCN (Wei et al., 2020) modify the structure of the interaction graph adaptively by discovering and pruning potential false-positive edges; and LATTICE (Zhang et al., 2021) mines the latent structure between items by learning an item-item graph. Graph convolutional operations are performed on both item-item and user-item interaction graphs to learn user and item representations.
In this paper, we present a simple yet effective GCN-based recommendation model that propagates embeddings over semantic aspect-based user-item interaction graphs. These graphs are built using the semantic aspect-aware interactions extracted by LLM from user reviews. Our proposed model outperforms existing GCN-based recommendation methods in terms of accuracy and interoperability.